OpenStudy (anonymous):

Convert the equation to the standard form for a hyperbola by completing the square on x and y. y^2 - 25x^2 + 4y + 50x - 46 = 0

OpenStudy (anonymous):

lol another multiple choice?

OpenStudy (anonymous):

Yes A. (x+2)^2/25 - (y-1)^2 = 1 B. (x-1)^2 - (x+2)^2/25 = 1 C. (x+2)^2/25 - (x-1)^2 = 1

OpenStudy (anonymous):

ok well that makes it much easier

OpenStudy (anonymous):

actually there must be something wrong here, because the \(y^2\) term comes first

OpenStudy (anonymous):

\[y^2 - 25x^2 + 4y + 50x - 46 = 0 \]

OpenStudy (anonymous):

these are my only options :/

OpenStudy (anonymous):

if the last one is really \[\frac{(y+2)^2}{25}-(x-1)^2=1\] then go with that one

OpenStudy (anonymous):

thank you, i'm going to go with it.